A simple random sample of 60 items resulted in a sample mean of 64. The population standard deviation is 12.

a. Compute the 95% confidence interval for the population mean (to 1 decimal).

(  ,   )

b. Assume that the same sample mean was obtained from a sample of 120 items. Provide a 95% confidence interval for the population mean (to 2 decimals).

(  ,   )
Answer #1

Solution: given that, n = 60 x=64 6= 12 a. At the z is 95% confidence level Xalgsl x=1-0.95 x=0.05 42 = 0.025 7042 = 70.025 =b) n = 120 is | At 95% confidence level the z dol-95 2=1-0.95 x=0.05 42 = 0.025 2x2= Zo.025= 1.96 margin of error E= Z42 X 6